# Segment Anything Model for Road Network Graph Extraction

### v1

@@ 4.6 Ablation studies
Table 4: The 7 SAM-Road variants compared for ablation studies. Opt: use optimized score thresholds. SAM: using pre-trained SAM. TFM: using a transformer for topology prediction. Offset: taking relative offsets in topology decoder. F-target: topology decoder takes target node feature. Itsc: predict intersection masks.
@@
How important is Using the pre-trained SAM model? A vs B proves it is critical. We repeated the experiment with the same ViT-B architecture with only ImageNet1K and MAE pre-training, and the results were far worse. This is not surprising, as City-scale and SpaceNet datasets are quite small in this era, especially when using large patch sizes (E.g. 512), giving it a sense of few-shot learning. The large-scale pre-training on datasets like SA-1B used by SAM seems critical for the generalization capability. Maybe it’s due to this reason that the baseline methods have to rely on smaller patches for more training examples and adopt weaker backbones with more inductive bias like CNNs.
@@
Whether using the learning-based topology decoder. A vs G shows that it’s critical for SAM-Road’s performance. Intuitively, a naive method that might achieve a similar effect is just to run a pathfinding algorithm between a pair of vertices, using the road existence map as the cost field, and see if there’s a sufficiently low-cost path between the two without passing through other vertices. We implemented such a variant G using an A-star algorithm. Metric is much worse, and qualitatively as shown in Figure  5. This approach can mess up intersections, overpasses, and close parallel roads.
Whether predicting the intersection vertices. This is answered by A vs F. Predicting intersection points is important for building correct intersection structures as shown in  5. Without it, both metrics drop.

### v2

@@ 4.6 Ablation Studies
Table 4: The SAM-Road variants compared for ablation studies. Opt: using optimized score thresholds. SAM: using pre-trained SAM. TFM: using a transformer for topology prediction. Offset: taking relative offsets in topology decoder. F-target: topology decoder takes target node feature. Itsc: predict intersection masks.
@@
How important is using the pre-trained SAM model? A vs B proves it is critical. We repeated the experiment with the same ViT-B architecture with only ImageNet1K and MAE pre-training [32], and the results were far worse. This is not surprising, as City-scale and SpaceNet datasets are quite small in this era, especially when using large patch sizes (E.g. 512), resembling few-shot learning. The large-scale pre-training on datasets like SA-1B used by SAM seems critical for the generalization capability. Maybe it’s due to this reason that the baseline methods have to rely on smaller patches for more training examples and adopt weaker backbones with more inductive bias like CNNs.
@@
Whether using the learning-based topology decoder. A vs G shows that it’s critical for SAM-Road’s performance. Intuitively, a naive method that might achieve a similar effect is just to run a pathfinding algorithm between a pair of vertices, using the road existence map as the cost field, and see if there’s a sufficiently low-cost path between the two without passing through other vertices. We implemented such a variant G using an A-star algorithm. Metric is much worse, and qualitatively as shown in Figure  6. This approach can mess up intersections, overpasses, and close parallel roads.
Whether predicting the intersection vertices. This is answered by A vs F. Predicting intersection points is important for building correct intersection structures as shown in  6. Without it, both metrics drop.